OpenStudy (anonymous):

HOW DO I DIVIDE THIS 1 1/3 yards by 6.5

OpenStudy (anonymous):

its the same as saying\[{11\over3} \div ~6{1\over2}~~~\implies~~~{11\over 3} \times {13\over 2}\implies~~~{11\over 3} \times {2\over 13}\] you change 6.5 into an improper fraction and dividing fractions is the same as multiplying by the reciprocal :)

OpenStudy (anonymous):

now you just solve\[{11\over3}\times {2\over 13}\] and thats your answer! :)

OpenStudy (anonymous):

@emelylove do you get it? :)

OpenStudy (anonymous):

yeah thanks i am your fan your awsome

OpenStudy (anonymous):

your welcome! anytime ;) ...thanks!! :D

OpenStudy (anonymous):

: )